OpenStudy (anonymous):

Convert log_{3}45 to natural logarithms then evaluate.

OpenStudy (anonymous):

\[\log_{3}45 \]

OpenStudy (anonymous):

Use this rule : \[\log_ab=\frac{\ln b}{\ln a}\]

OpenStudy (anonymous):

And you can remark that : \[45=3^2\times 5\]

OpenStudy (anonymous):

ln45/ln3?

OpenStudy (anonymous):

@ducksonquack yes : \[\log_345=\frac{\ln45}{\ln3}=\frac{\ln(3^2\times 5)}{\ln 3}=\cdots\]

OpenStudy (anonymous):

so 3.5?

OpenStudy (anonymous):

Yes, by using the calculator ;)

OpenStudy (anonymous):

Thanks!